Proceedings of the Ninth Prague Topological Symposium Contributed papers from the symposium held in Prague, Czech Republic, August 19–25, 2001

CHAINABLE SUBCONTINUA

EDWIN DUDA

Abstract. This paper is concerned with conditions under which a metric continuum (a compact connected metric space) contains a nondegenerate chainable continuum.

This paper is concerned with conditions under which a metric continuum (a compact connected metric space) contains a non-degenerate chainable continuum.

By R.H. Bing’s theorem eleven [2] if a metric continuum X contains a non-degenerate subcontinuum H which is hereditarily decomposable, hereditarily unicoherent, and atriodic, then H is chainable.

The following papers give examples of continua with the property that each non-degenerate subcontinuum is not chainable. G.T. Whyburn [16]. R.D. Anderson and G. Choquet [1]. A. Lelek [6] gives an example of a planar weakly chainable continuum each non-degenerate subcontinuum of which separates the plane and thus contains no non-degenerate chainable subcontinuum. W.T. Ingram [5] gives an example of an hereditarily indecomposable tree-like continuum such that each non-degenerate subcontinuum has positive span and hence is not chainable.

C.E. Burgess in [3] shows if a continuum M is almost chainable and K is a proper subcontinuum of M which contains an endpoint p of M, then K is linearly chainable with p as an end point. A continuum M is almost chainable if, for every positive number ε, there exists an ε-covering G of M and a linear chain C(L_{1}, L_{2}, …, L_{n}) of elements of G such that no L_{i} (1 ≤ i < n) intersects an element of G − C and every point of M is within a distance ε of some element of C. He also shows if M is almost chainable, then M is not a triod and M is unicoherent and irreducible between some two points. Examples show M can contain a triod or a non-unicoherent subcontinuum.

2000 Mathematics Subject Classification. 54F20.

Key words and phrases. chainable continuum.

Edwin Duda, Chainable subcontinua, Proceedings of the Ninth Prague Topological Symposium, (Prague, 2001), pp. 71–73, Topology Atlas, Toronto, 2002; arXiv:math.GN/0204122.

If X and Y are metric continua and if X can be ε-mapped onto Y for all positive ε and Y has a non-degenerate chainable continuum then so does X. This result suggests considering inverse limit spaces. At this stage we refer to a result from the paper of S. Mardes̆ić and J. Segal [9] Theorem 1, p. 148: “Every π-like continuum X is the inverse limit of an inverse sequence {P_{i}; π_{ij}} with bonding maps π_{ij} onto and with polyhedra P_{i} ∈ π. A continuum is π-like if it can be ε-mapped onto some polyhedron in π for each positive ε. E. Duda and P. Krupski [4] showed that a k-junctioned metric continuum, k a non-negative integer, has at most k points such that any continuum which contains none of the k points is chainable. A metric continuum is said to be k-junctioned if it is the inverse limit of graphs each of which has at most k branch points, with surjective bonding maps. A continuum is called finitely junctioned if it is k-junctioned for some non negative integer k.

Suppose now X is a tree-like continuum. Then for each ε > 0 X can be mapped onto a tree. By a result quoted above X is the inverse limit of a sequence of trees with surjective bonding maps. X = lim_{⟵}{T_{n}, f_{nm}}. Let f_{n} : X → T_{n} be the standard projection map and let

P_{n} = U{f_{n}^{−1}(q)|q is a branch point}.

Since T_{n} has at most a finite number of branch points (points of order ≥ 2) P_{n} is closed in X. If the union of the P_{n} is not dense in X then X contains a non-degenerate chainable continuum. Actually it is sufficient that {P_{n}} have a subsequence whose union is not dense in X.

Lets now consider a non-degenerate metric continuum in X with span equal to zero. The notion of span was defined by A. Lelek [7]. In the paper [8] he showed continua with span zero are atriodic and tree-like.

There is a series of papers by L.G. Oversteegen and E.D. Tymchatyn which develop properties of spaces with spans equal to zero or sufficient conditions that a space have a span equal to zero [12, 15, 13, 14]. Also by L.G. Oversteegen [11].

It is interesting to note that a chainable continuum X can be ε-mapped onto any fixed dendrite. Thus for any tree T, by the result of Mardes̆ić and Segal quoted above, X is the inverse of a sequence of T ’s.

In the paper [10] P. Minc shows an inverse limit of trees with simplicial bonding maps having surjective span zero is chainable.

References

[1] R. D. Anderson and Gustave Choquet, A plane continuum no two of whose nondegenerate subcontinua are homeomorphic: An application of inverse limits, Proc. Amer. Math. Soc. 10 (1959), 347–353. MR 21 #3819

[2] R. H. Bing, Snake-like continua, Duke Math. J. 18 (1951), 653–663. MR 13,265a [3] C. E. Burgess, Homogeneous continua which are almost chainable, Canad. J. Math.

13 (1961), 519–528. MR 23 #A3551

[4] Edwin Duda and Paweł Krupski, A characterization of finitely junctioned continua, Proc. Amer. Math. Soc. 116 (1992), no. 3, 839–841. MR 93a:54031 [5] W. T. Ingram, Hereditarily indecomposable tree-like continua. II, Fund. Math. 111

(1981), no. 2, 95–106. MR 82k:54056

[6] A. Lelek, On weakly chainable continua, Fund. Math. 51 (1962/1963), 271–282. MR 26 #742

[7] , Disjoint mappings and the span of spaces, Fund. Math. 55 (1964), 199–214. MR 31 #4009

[8] , The span of mappings and spaces, Topology Proc. 4 (1979), no. 2, 631–633. MR 82c:54031

[9] Sibe Mardešić and Jack Segal, ε-mappings onto polyhedra, Trans. Amer. Math. Soc. 109 (1963), 146–164. MR 28 #1592

[10] Piotr Minc, On simplicial maps and chainable continua, Topology Appl. 57 (1994), no. 1, 1–21. MR 95c:54056

[11] Lex G. Oversteegen, On span and chainability of continua, Houston J. Math. 15 (1989), no. 4, 573–593. MR 91g:54049

[12] Lex G. Oversteegen and E. D. Tymchatyn, Plane strips and the span of continua. I, Houston J. Math. 8 (1982), no. 1, 129–142. MR 84h:54030

[13] , On the span of weakly-chainable continua, Fund. Math. 119 (1983), no. 2, 151–156. MR 85j:54051

[14] , On span and weakly chainable continua, Fund. Math. 122 (1984), no. 2, 159–174. MR 85m:54034

[15] , Plane strips and the span of continua. II, Houston J. Math. 10 (1984), no. 2, 255–266. MR 86a:54042

[16] G. T. Whyburn, A continuum every subcontinuum of which separate the plane, Amer. J. Math. 52 (1930), 319–330.

University of Miami, Department of Mathematics, PO Box 249085, Coral Gables, FL 33124-4250

E-mail address: e.duda@math.miami.edu